OpenStudy (anonymous):

Problem: For each of the following equations, find the value(s) of the constant α so that the equation has exactly one solution, and determine the solution for each value. (a) αx2 + x + 1 = 0 (b) x2 + αx + 1 = 0 (c) x2 + x + α = 0 (d) x2 + αx + 4α + 1 = 0 Need Help PLEASE.. THANKS IN ADVANCE..

OpenStudy (abb0t):

Factor each out into two equations so you have (a+b)(a+c) and set them equal to zero to solve for your "unknown"

OpenStudy (swissgirl):

Well I would suggest a different method especially for a

OpenStudy (anonymous):

I would suggest the quadratic formula.

OpenStudy (anonymous):

Then make it so \(\alpha \) gives you repeated roots.

OpenStudy (anonymous):

You basically want it so that: \[ \sqrt{b^2-4ac} = 0 \]Where \(a\), \(b\), and \(c\) are coefficients: \[ ax^2+bx+c \] @rashley284 Make sense?

OpenStudy (swissgirl):

YAAAAA thats it

OpenStudy (anonymous):

thank you @wio & @swissgirl

OpenStudy (anonymous):

\[ x = \frac{-b\pm \sqrt{b^2-4ac}}{2a} \]The \(\pm \) will be singular if followed by a \(0\)... if you wanna understand the reasoning.

OpenStudy (swissgirl):

We are looking for one solution which means there will only be one root. When the discriminant is 0 there is only one root like @wio has shown